# LIVS: A Pluralistic Alignment Dataset for Inclusive Public Spaces

### v1

@@ Results.
Out of 2,200 new comparisons, annotators chose the DPO-aligned model in 700 (32%) instances and the baseline in 300 (14%), marking the remaining 1,100 (50%) as neutral (See Appendix I). Criteria with more training annotations (e.g., Comfort, Invitingness) showed stronger improvements under DPO, whereas Inclusivity and Diversity had higher neutral ratings. Qualitative inspections indicated that DPO led to clearer walkways and seating configurations but did not consistently render detailed features (e.g., ramps, tactile features, multilingual signage) (Figure 5). Overall, the results suggest that multi-criteria DPO can align T2I outputs to local preferences while retaining considerable variability across criteria (see Appendix G for additional details).
@@
Comparing Figure 4 (overall dataset) with Figure 5 (evaluation set) suggests a similar distribution of annotations and neutral responses: criteria with more training annotations (e.g., Comfort, Invitingness) generally exhibit stronger DPO preference in the evaluation. This pattern indicates that additional data may further improve multi-criteria alignment through DPO. However, we currently do not leverage neutral labels, and majority voting in multi-criterion comparisons may overlook nuanced disagreements across criteria. Future work could explore methods for incorporating neutral ratings and resolving multi-label conflicts without collapsing them into binary outcomes. Further analysis of criterion-level variance is provided in the Appendix I.
@@
Figure 8: Rates of neutral annotations for human-written (Methods 0) versus GPT-4o-generated prompts (Methods 1-4). Method 0 indicates that human-authored prompts elicited more decisive preferences (fewer neutral annotations), suggesting clearer visual distinctions
@@
Human-authored prompts led to fewer neutral ratings, suggesting they produced more visually distinct outcomes (Figure 8). GPT-4o-generated prompts exhibited higher neutrality, possibly due to less contextual specificity. These findings indicate that prompt design can influence annotators’ perceptions of alignment differences.
@@
We observed that individuals from various intersectional groups (e.g., participants with mobility challenges) typically assigned lower Accessibility or Safety scores. This variation underscores the need for local, intersectional feedback in T2I alignment for urban planning, as a single global objective cannot capture such diverse preferences. See Appendix Section H.1, Figures 15 and  16 for further details.
@@ 5 Implications
Our findings contribute to ongoing discourse on trustworthy and multi-objective machine learning. Below we highlight several implications:
Multi-Criteria Overlaps and Conflicts. Some participants valued certain criteria (like Diversity) above others (such as Comfort). Others perceived Safety and Comfort as interlinked, finding it hard to separate physical hazards from environmental conditions. Future alignment methods might benefit from hierarchical or weighted objective formulations, capturing partial dependencies among criteria (Gabriel & Ghazavi, 2021; Stray, 2020; Li et al., 2024; Sorensen et al., 2024).
Neutral Annotations as a Signal. Approximately half of the final evaluations were rated as neutral, with participants commenting that neither image fully captured local expectations. Rather than viewing these neutral responses as alignment failures, we interpret them as indications that the alignment adjustments were either too subtle or insufficient to reflect the complexity of participant feedback. Prompt variations also influenced these outcomes, as some prompts elicited more ambiguous or less distinct visual differences. Taken together, neutral ratings highlight the importance of richer, context-specific alignment methods that account for both the prompt and the underlying multi-criteria data, including contradictory user needs, and underscore the need for further research (Hosking et al., 2024; Sorensen et al., 2024).
@@ 6 Conclusion
Availability. The LIVS dataset, including citizens’ self-identification markers (with consent), will be made available for research purposes. This release aims to establish a benchmark for pluralistic alignment in text-to-image generation and supports both criterion-specific and user-specific customization. Future work can leverage these granular annotations to develop personalized models and adaptive fine-tuning strategies that more effectively address the unique needs of diverse user groups.
@@ Appendix G Qualitative Observations
Figures 12, 13, and 14 compare baseline SDXL images against versions finetuned on specific or multiple LIVS criteria. While DPO alignment often improves features such as smooth surfaces or clearer paths, certain elements (e.g., ramps, multilingual signage) appear inconsistently, indicating the model’s limited capacity for rendering specialized details.
@@ H.1 Case Study IV: Scoring Patterns of SDXL Images by Intersectional Identities
Figures 15 and 16 plot average raw scores from intersectional groups (e.g., disability status $\times$ race/ethnicity) for SDXL-generated images across the six LIVS criteria. Participants with mobility challenges generally assigned lower Accessibility and Safety scores, highlighting the importance of soliciting localized, intersectional feedback in T2I alignment for public-space design.

### v2

@@ Results.
Out of 2,200 new comparisons, annotators chose the DPO-aligned model in 700 (32%) instances and the baseline in 300 (14%), marking the remaining 1,100 (50%) as neutral (See Appendix I). Criteria with more training annotations (e.g., Comfort, Invitingness) showed stronger improvements under DPO, whereas Inclusivity and Diversity had higher neutral ratings (Figure 6). Qualitative inspection revealed that DPO produced clearer walkways and seating configurations but did not consistently generate detailed features (e.g., ramps, tactile paving, multilingual signage). Examples of baseline, single-criterion, and multi-criteria fine-tuned outputs are shown in Figure 7. Overall, the results suggest that DPO improves the alignment of T2I outputs with local preferences but continues to exhibit substantial variability across criteria. The high proportion of neutral ratings indicates that DPO alone is insufficient to capture the complexity of community preferences, underscoring the need for more nuanced and adaptive alignment algorithms (see Appendix G for additional details).
@@
Comparing Figure 4 (overall dataset) with Figure 5 (evaluation set) suggests a similar distribution of annotations and neutral responses: criteria with more training annotations (e.g., Comfort, Invitingness) generally exhibit stronger DPO preference in the evaluation. This pattern indicates that additional data may further improve multi-criteria alignment through DPO. However, we currently do not leverage neutral labels, and majority voting in multi-criterion comparisons may overlook nuanced disagreements across criteria. Specifically, when annotators provide conflicting preferences across the three assigned criteria, the final label is determined by majority vote—e.g., two preferences for the left image and one for the right results in a binary label favoring the left. This simplification disregards the internal variation across criteria and may obscure which dimensions drive preference. Future work could explore methods for incorporating neutral ratings and resolving multi-label conflicts without collapsing them into binary outcomes. Further analysis of criterion-level variance is provided in Appendix I.
@@
Figure 6: Average DPO preference by criterion. Bars show the mean tendency toward DPO-aligned outputs. Performance tends to improve in criteria with higher annotation counts.
@@
Human-authored prompts (Method 0) led to fewer neutral ratings, suggesting they produced more visually distinct outcomes (Figure 9). In contrast, GPT-4o-generated prompts (Methods 1–4) exhibited higher rates of neutrality, possibly due to reduced contextual specificity. As shown in Figure 9, the proportions of neutral annotations vary across methods, indicating that prompt design significantly influences annotators’ perceptions of alignment differences.
@@
We observed that individuals from various intersectional groups (e.g., participants with mobility challenges) typically assigned lower Accessibility or Safety scores. This variation underscores the need for local, intersectional feedback in T2I alignment for urban planning, as a single global objective cannot capture such diverse preferences. See Appendix Section H.1, Figures 16 and  17 for further details.
Figure 9: Proportion of neutral annotations for human-written (Methods 0) versus GPT-4o-generated prompts (Methods 1-4). Method 0 indicates that human-authored prompts elicited more decisive preferences (fewer neutral annotations), suggesting clearer visual distinctions
@@ 5 Implications
Our findings inform pluralism and multi‐objective machine learning. Implications follow:
Multi-Criteria Overlaps and Conflicts. Some participants valued certain criteria (like Diversity) above others (such as Comfort). Others perceived Safety and Comfort as interlinked, finding it hard to separate physical hazards from environmental conditions. Future alignment methods might benefit from hierarchical, weighted, context-sensitive, or dynamic objective formulations, capturing partial dependencies among criteria (Gabriel & Ghazavi, 2021; Stray, 2020; Li et al., 2024; Sorensen et al., 2024).
Neutral Annotations as a Signal. Approximately half of the final evaluations were rated as neutral. While these outcomes may initially suggest a lack of meaningful improvement, they may also reflect genuinely balanced preferences or participants’ difficulty in discerning visual cues for certain criteria—particularly Inclusivity and Diversity. In several interviews, participants noted that subtle or symbolic elements were not always visually salient. Rather than interpreting neutral responses as alignment failures, we consider them indicative of alignment adjustments that were either too subtle or inadequate to capture the complexity of participant feedback. Prompt variations also contributed to this ambiguity, as some prompts yielded less distinct visual differences. Overall, neutral ratings emphasize the importance of richer, context-sensitive alignment methods that consider both prompt content and multi-criteria data, including conflicting user needs, and highlight the need for further research (Hosking et al., 2024; Sorensen et al., 2024).
Methodological Constraints and Alignment Strategy. Because DPO requires binary labels, we collapsed multi-criteria feedback into single preference signals during training. This simplification overlooks intra-sample disagreement and limits the model’s capacity to capture the full complexity of pluralistic preferences. Future alignment strategies should accommodate multi-dimensional annotations directly, enabling more nuanced modeling of diverse community feedback.
@@ 6 Conclusion
Availability. The LIVS dataset—including citizen-provided self-identification markers (with consent)—is available for research purposes at mid-space.one. This release aims to establish a benchmark for pluralistic alignment in text-to-image generation and supports both criterion-specific and user-specific customization. Future work can leverage these granular annotations to develop personalized models and adaptive fine-tuning strategies that more effectively address the unique needs of diverse user groups.
@@
Acknowledgements
We thank Emmanuel Beaudry-Marchand for developing the foundational schema of the Aipithet Platform used during the annotation process. We also acknowledge Toumadher Ammar for her assistance with the workshops, and Jerome Solis for his support with coordination. We are grateful to Sarah Tannir, Leandry Jieutsa, Adèle Kremer, Frédérique Roy, and Roxane Kasprzyk for their contributions.
We appreciate the collaboration of our partner organizations and architectural offices, including Enclume, Sid Lee Architecture, Dark Matter Labs, IVADO, and the Canadian Commission for UNESCO.
This research was funded by the Soutien aux initiatives avec les collectivités et les entreprises – Collaboration avec les organismes communautaires program of the Université de Montréal, and the Bourse IA ESP. Additional support was provided by Mitacs and the Fonds de recherche du Québec – Société et culture (FRQSC) Doctoral Scholarship.
We thank all 30 community organizations in Montreal that supported citizen collaboration. The following were among the most engaged and actively involved: the Congolese Community Center of Montreal, Altergo, La Maisonnée, Cummings Centre, Projet Changement, Women’s Center of Plateau Mont-Royal, LGBTQ+ Community Center of Montreal, Marguerite-Bourgeoys Hub, RÉZO, Afrique au Féminin, L’Agence On est là!, and the Montreal Women’s Groups Table.
This work was enabled in part by computing resources provided by Mila.
@@ Appendix G Qualitative Observations
Figures 13, 14, and 15 compare baseline SDXL images against versions finetuned on specific or multiple LIVS criteria. While DPO alignment often improves features such as smooth surfaces or clearer paths, certain elements (e.g., ramps, multilingual signage) appear inconsistently, indicating the model’s limited capacity for rendering specialized details.
@@ H.1 Case Study IV: Scoring Patterns of SDXL Images by Intersectional Identities
Figures 16 and 17 plot average raw scores from intersectional groups (e.g., disability status $\times$ race/ethnicity) for SDXL-generated images across the six LIVS criteria. Participants with mobility challenges generally assigned lower Accessibility and Safety scores, highlighting the importance of soliciting localized, intersectional feedback in T2I alignment for public-space design.